A square support unit in a television is made with a side measuring 4 centerimeters. A new model being designed for the upcoming year will have a larger square with a side measuring 4.3 centimeters. By how much will the area of the square be increased?
Answer by rfer(10432)

4*4=16
4.3*4.3=18.49
18.49-16=2.49cm increase